Ticket Number: HEA-NOC/20080131-6 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20080131 16:37 UTC Problem Start: 20080205 10:00 UTC
Ticket Update: 20080206 10:17 UTC Problem End: 20080206 10:00 UTC
UCC is to be moved from their current connection to our Cork access router
to instead route to the new CRS routers via Bluenet.
While the majority of the configuration changes will be without impact, the
actual transition from the direct connection to the Cork PoP to the new
topology will cause an extremely brief outage of far less than one minute.
The UCC connection should be considered at risk during the entire
20080131 16:30 BN Opened change request, sent initial configuration draft
20080201 16:07 BN Confirmed dates with UCC. The non-service affecting
work will take place on Tuesday 5th of Feb. The work requiring a brief
outage will take place on Wednesday 6th between 08:00 and 09:00.
20080206 10:16 BN Work completed as expected within the window. No
noticeable outage recorded, although routing was briefly asynchronous.
New connections will be monitored and ticket kept open for at least
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20080131-6
HEAnet Limited. Registered in Ireland, No. 275301.